divide 184 in 2 parts such one-third may exceed one-seventh of another part by 8
P.S= PLS GIVE THE ANSWER IN LINEAR EQUATION

Answer by rothauserc(4718)

we have two equations in two unknowns, let x be one part and y be the other part
:
1) x + y = 184
2) (x/3) = (y/7) + 8
:
solve 1) for x and substitute for x in 2)
:
x = 184 - y
( (184 - y) / 3 ) = (y/7) + 8
:
multiply both sides of = by 21
:
7(184-y) = 3y + 168
:
1288 -7y = 3y + 168
:
10y = 1120
:
*******
y = 112
x = 72
*******
:
to check this substitute for x and y in 2)
:
72/3 = 112/7 + 8
:
24 = 16 + 8
:
24 = 24
:
answer checks :-)
: